Your uncle walks in, jingling the coins in his pocket. He grins at you and tells you that you can have all the coins if you can figure out how many of each kind of coin he is carrying. The twenty-six coins in his pocket are all dollars and quarters, and they add up to seventeen dollars in value. How many of each coin does he have?
Answer by mananth(16946)

let $1 be x
quarters will be 26-x
...
25(26-x)+100x=1700
650-25x+100x=1700
75x=1700-650
75x=1050
x=1050/75
x=14 dollar coins
12 quarters
...
m.ananth@hotmail.ca